OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

jimthompson5910 (jim_thompson5910):

Hint: −5|x + 1| = 10 −5|x + 1|/(-5) = 10/(-5) ... Divide both sides by -5. |x+1| = -2

jimthompson5910 (jim_thompson5910):

now you need to find the solutions to |x+1| = -2

OpenStudy (anonymous):

x = −3 and x = 1

OpenStudy (anonymous):

or x = −1 and x = 3

jimthompson5910 (jim_thompson5910):

are there any solutions to |x| = -2

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

no solution

jimthompson5910 (jim_thompson5910):

good, if you have some absolute value equal to a negative number, then there are no solutions